Familiar with the rules for connecting the conductors, Jurgita decided to experiment with the garland of the Christmas tree. Determine how the luminosity of the lantern bulbs changed when Jurgita turned on one additional bulb; took out one bulb and plugged the contacts in place.
Answer #1

In this question we have to find the effect on the luminosity of the bulbs when an additional bulb is turned ON and also when a bulb is removed from the arrangement.Christmas tree. Let Jurgita hap n bulbs in her Let Supply Voltage be v and Resistance g each bulb . be R. en bulbs TAAR AAN Abulb is turned on. Now when n additional Then I = V (n+1) R power consumed by each Then . Pi= I?R - Pi= V² (n+1²R bulb be P1from the when one bulb is removed gasland q Christmas tree. Remaining bulbs = (n-1) .: Cuorent I = V (n-1) R power in this ca